OpenStudy (anonymous):

Use order of operations to simplify. 6/5 divided by (1/5 + 1/15) = (type a mixed number)

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Is this the question? \( \dfrac{6}{5} \div \left( \dfrac{1}{5} + \dfrac{1}{15} \right) = \)

OpenStudy (tkhunny):

Are you going to show us some of your work? Go ahead and add those two in the denominator.

OpenStudy (anonymous):

First you do the equation in the parentheses. = 1/5 + 1/15 = ((1 × 15) + (1 × 5)) / (5 × 15) = (15 + 5) / 75 = 20/75 = 4/15 and now divide 6/5 by 4/15 = 6/5 ÷ 4/15 = (6 × 15) / (4 × 5) = 90/20 = 9/2 = 4 1/2